i want to know how youh turn decimals into percents | or the other way around.. or turn fractions into percents/ decimals ?
Answer by rfer(12726)

1/4=.25=25%
1/2=.5=50%
2/5=.40=40%
Divide a fraction, the bottom into the top to get a decimal.
Move the decimal two places to the right to change a decimal into %
With a % move the decimail two places to the left for a decimal.
decimal to a fraction if two place put the number over 100.
.40=40/100=reduce 4/10=2/5
.5=5/10=1/2